OpenStudy (anonymous):

(t-3)7=6t+21 show work

OpenStudy (valpey):

The answer to the universe and everything.

OpenStudy (anonymous):

come on man, you know you have to distibute that 7 into the parentheses

OpenStudy (chaise):

(t-3)7=6t+21 7t-21=6t+21 7t=6y+42 t = 42